# Bimanual Manipulation of Steady Hand Eye Robots with Adaptive Sclera Force Control: Cooperative vs. Teleoperation Strategies

### v1

@@ Abstract
Performing intricate eye microsurgery, such as retinal vein cannulation (RVC), as a potential treatment for retinal vein occlusion (RVO), without the assistance of a surgical robotic system is very challenging to do safely. The main limitation has to do with the physiological hand tremor of surgeons. Robot-assisted eye surgery technology may resolve the problems of hand tremors and fatigue and improve the safety and precision of RVC. The Steady-Hand Eye Robot (SHER) is an admittance-based robotic system that can filter out hand tremors and enables ophthalmologists to manipulate a surgical instrument inside the eye cooperatively. However, the admittance-based cooperative control mode does not address crucial safety considerations, such as minimizing contact force between the surgical instrument and the sclera surface to prevent tissue damage. An adaptive sclera force control algorithm was proposed to address this limitation using an FBG-based force-sensing tool to measure and minimize the tool-sclera interaction force. Additionally, features like haptic feedback or hand motion scaling, which can improve the safety and precision of surgery, require a teleoperation control framework. A single-manual adaptive teleoperation control mode was implemented on SHER 2.1 in previous work for the first time in robot-assisted retinal microsurgery. We extended the previous work into a bimanual adaptive teleoperation (BMAT) control mode using SHER 2.0 and SHER 2.1 and compared its performance with a bimanual adaptive cooperative (BMAC) mode. Both BMAT and BMAC modes were tested in sitting and standing postures during a vessel-following experiment under a surgical microscope. It is shown, for the first time to the best of our knowledge in robot-assisted retinal surgery, that integrating the adaptive sclera force control algorithm with the bimanual teleoperation framework enables surgeons to safely perform bimanual telemanipulation of the eye without over-stretching it, even in the absence of registration between the two robots. Experimental results demonstrate the effectiveness of the proposed BMAT control framework.
@@ I INTRODUCTION
Retinal vein occlusion (RVO) has been identified as the second most prevalent potentially blinding retinal vascular disease, impacting an estimated 28 million individuals globally, according to a meta-analysis study [1]. This condition stems from the occlusion of a retinal vein, leading to eye problems including vision loss. Presently, there is no established standard for directly treating RVO with a surgical approach. Retinal vein cannulation (RVC) has been attempted in order to deliver a therapeutic agent directly into the affected vein [2], however, its applicability is highly challenging due to the micron scale of retinal veins and human hand tremor [3]. The problem of scale is presently best addressed with the assistance of a surgical robotic system. Of note, the diameter of the largest retinal veins is on the order of $150\mu$m [4], and the root mean square (RMS) amplitude of hand tremor for an ophthalmic surgeon is measured to be $182\mu$m [5].
@@
Several surgical robotic systems have been developed to filter out physiological hand tremors and improve the safety of retinal surgery. Some examples of these systems include the Steady Hand Eye Robot (SHER) [6, 7], Preceyes [8], to name a few [9, 10, 11].
@@
Figure 1: Bimanual teleoperation architecture with adaptive force control algorithm. The system consists of two Phantom Omni robots, the SHER 2.0 and SHER 2.1, a surgical microscope, two FBG-based force-sensing surgical instruments attached to the SHERs’ end-effector to measure the tip force and sclera force, a monitor, and two foot pedals to activate and control the SHERs impedance. The figure shows the high-level adaptive force controller, a mid-level optimizer, and a low-level Galil joint velocity controller for SHER 2.0 and SHER 2.1.
@@
The SHER 2.0 and SHER 2.1 are two robot versions developed at Johns Hopkins University. The main control algorithm implemented on SHERs is an admittance-based control where the user and the robot cooperatively manipulate a surgical tool. Using a force/torque sensor attached to the robot handle, the exerted force by the user is measured and converted to a desired end-effector velocity to control the robot [7]. An important limitation of this control algorithm is that the force/torque sensor is not sensitive enough to measure the minimal interaction force existing on the surgical instrument at the sclera entry (RCM), which could result in applying excessive force to the sclera distorting and potentially causing damaging. To overcome this issue, Ebrahimi et al. proposed an adaptive sclera force control algorithm [12] using an FBG-based force-sensing tool to automatically minimize the sclera force and keep it within a suggested safe range of 120 mN [13]. Notably, FBG sensors have applications in flexible needles or continuum robots, serving purposes such as shape sensing [14] or force sensing [15]. A comparison of the adaptive force control algorithm and the admittance-based control algorithm performance was conducted using a vessel-following experiment. An eye phantom was used where the SHER 2.1 held the primary force-sensing surgical tool and cooperatively manipulated the user’s dominant hand. Simultaneously, a conventional surgical instrument was directly held by the non-dominant hand, serving as a secondary tool for re-orienting the eye under a surgical microscope [12]. Despite reducing the applied force to the sclera by the dominant hand compared to the admittance-based control mode, controlling the force magnitude on the secondary hand-held tool is impossible, and applying excessively unsafe force to the sclera cannot be avoided. Also, this evaluation took place within a single-manual cooperative control mode without teleoperation capability. Integrating this adaptive force control algorithm in a teleoperation modality has the potential to offer surgeons advanced capabilities, contributing to an improvement in patient safety. Esfandiari et al. implemented this algorithm in a teleoperation modality and compared its performance with a cooperative control mode in a single-manual framework [16].
@@
In this work, we extended the previous single-manual teleoperation framework to a bimanual adaptive teleoperation control framework integrated with an adaptive sclera force control algorithm on the SHER 2.0 and the SHER 2.1 systems using two force-sensing surgical instruments. This framework enables surgeons to safely conduct bi-manual teleoperation on the eye, preventing any stretching of the eye, even without registering the two robots. For the first time in robot-assisted retinal microsurgery applications, we compared the performance of the proposed bimanual adaptive teleoperation control mode and a bimanual adaptive cooperative control mode integrated with the same adaptive force control algorithm. This evaluation occurred during a vessel-following experiment on an eye phantom under a surgical microscope for both sitting and standing postures for each control mode. The objective was to assess and compare the performance of the BMAT and BMAC control modes and also the effect of user’s posture on their performance.
The two FBG-based force-sensing surgical instruments are calibrated using a method outlined by [15]. This calibration enabled the estimation of the tip force, sclera force, and insertion depth by measuring the wavelength shifts of nine FBG sensors attached to each tool.
@@
The remaining sections of this paper are organized as follows. Methods are discussed in Section II. Section III describes the experimental setup and the experiment procedure. The experimental results are analyzed and discussed in Section IV. Finally, a conclusion is made in Section V.
Figure 2: Experiment setup includes the SHER 2.1 and the SHER 2.0, two PHANTOM Omni robots, two ATI force sensors, two force-sensing tools (attached to the SHERs’ handles) equipped with FBG sensors, two FBG interrogators, an eye phantom with four colored vessels and four target points (pins) attached to the end of each colored vessel, an armrest, and a surgical microscope.
@@ II Methods
The objective of this experiment is to extend a previously developed adaptive teleoperation control mode [16] to two eye robots, the SHER 2.0 and SHER 2.1, and to assess the performance of the two control frameworks: bimanual adaptive cooperative mode and bimanual adaptive teleoperation mode (see Fig. 1), in both sitting and standing postures. In both control modes, the sclera force magnitude is automatically controlled (without the user’s commands) by an adaptive force control algorithm [12] to avoid stretching and damaging the sclera. Following the suggestion of our clinical lead, we conducted a "vessel-following" experiment, a prototypical task in retinal surgery [17]. The experimental setup prepared to perform the vessel-following experiment provides a measurement and recording of SHERs’ kinematic information, the user’s hand force/torque during interaction with the SHERs’ handle, and the force applied to the surgical tool at the sclera point and the tooltip.
II-A Adaptive Force Control of the SHER
[…]
@@ II-A Adaptive Force Control of the SHER
The desired linear velocities of the SHER’s end-effector along the $x$ and $y$ directions in the body frame $\{B\}$ can be determined by the adaptive force controller using Equations (3) and (4):
@@
where $\Delta{f}_{i}=F_{si}-f_{di}$ denotes the sclera force error, and $\Gamma_{i}$ and $K_{fi}$ represent constant gains for the adaptation law and the force tracking error, respectively. $\alpha_{i}$ stands for the estimated tissue compliance, and $\dot{f}_{di}$ is the derivative of $f_{di}$. This adaptive sclera force control algorithm applies to both cooperative and teleoperation control modalities. In the cooperative control mode, the remaining components of the SHER end-effector velocity are generated by the admittance-based controller, whereas in teleoperation control mode, they are derived from the PHANTOM Omni’s end-effector velocity $\dot{X}_{o}^{b}$ (mapped to the SHER body coordinate $\{B\}$). The proposed teleoperation control mode, integrated with an adaptive sclera force control algorithm, consists of three primary components: a high-level controller, a mid-level optimizer with joint constraints, and a low-level joint velocity controller (Galil).
The high-level controller consists of two primary components: Firstly, there is a kinematic controller that translates the end-effector velocity of the Omni, denoted as $\dot{X}_{o}$, into the SHER’s body coordinate as $\dot{X}_{o}^{b}$. It is then pre-multiplied by $K=diag(\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4},\kappa_{5},\kappa_{6})$, a user-defined motion scaling matrix, to provide an intuitive motion on the SHER end-effector velocity, i.e., $\dot{X}_{des}^{b}=K\dot{X}_{o}^{b}$. Secondly, an adaptive sclera force controller that receives input from the sclera force components $F_{sx}$ and $F_{sy}$, which are measured in the body coordinate $\{B\}$ by the force-sensing instrument. This controller generates the desired adaptive end-effector velocity in the SHER body coordinate, denoted as $\dot{X}_{A}^{b}$ in Equation (3).
Figure 3: Vessel-following experiment procedure for four combinations of control modes and user posture including (a) standing BMAC, (b) standing BMAT, (c) sitting BMAC, and (d) sitting BMAT. (e) The user activates each robot using a foot pedal and follows a random order of colored vessels by the dominant (right) hand and touches a target point at the end of each vessel by the tooltip of the dominant hand and re-orients the eye phantom using the non-dominant hand under a surgical microscope.
This hybrid kinematic-force control method determines the desired SHER velocity by switching between $\dot{X}_{A}^{b}$ and $\dot{X}_{o}^{b}$ based on the magnitude of the sclera force. When the scleral force components fall below a safe threshold $T_{s}$, the SHER is controlled based on kinematic teleoperation using the Omni velocity $\dot{X}_{o}^{b}$; otherwise, the adaptive force controller takes charge, enforcing the SHER’s end-effector velocity through $\dot{X}_{A}^{b}$ to minimize scleral force components. As a result, the desired end-effector velocity of SHER, $\dot{X}_{des}^{b}$, is generated based on the following hybrid kinematic-force control algorithm:
@@
$\begin{bmatrix}(1-\Delta_{x})\kappa_{1}&0&\textbf{0}_{1\times 4}\\ 0&(1-\Delta_{y})\kappa_{2}&\textbf{0}_{1\times 4}\\ \textbf{0}_{4\times 2}&&diag(\kappa_{3},\kappa_{4},\kappa_{5},\kappa_{6})\end{bmatrix}\dot{X}_{o}^{b}$ (8)
@@
in which $\Delta_{i}$ ($i=x$ or $y$) is a binary variable automatically set by the control switching policy (Algorithm 1) based on the magnitude of sclera force components $F_{si}$ ($i=x$ or $y$). $\Delta_{i}=1$ means the adaptive force control algorithm for the corresponding sclera force component ($x$ or $y$) is activated. Otherwise, when $\Delta_{i}=0$, the adaptive force control algorithm gets deactivated, and Equation (14) becomes equivalent to the kinematic control $\dot{X}_{des}^{b}=K\dot{X}_{o}^{b}$.
Algorithm 1 Hybrid Kinematic-Force Control Switching Policy

### v2

@@ Abstract
Performing retinal vein cannulation (RVC) as a potential treatment for retinal vein occlusion (RVO) without the assistance of a surgical robotic system is very challenging to do safely. The main limitation is the physiological hand tremor of surgeons. Robot-assisted eye surgery technology may resolve the problems of hand tremors and fatigue and improve the safety and precision of RVC. The Steady-Hand Eye Robot (SHER) is an admittance-based robotic system that can filter out hand tremors and enables ophthalmologists to manipulate a surgical instrument inside the eye cooperatively. However, the admittance-based cooperative control mode does not safely minimize the contact force between the surgical instrument and the sclera to prevent tissue damage. Additionally, features like haptic feedback or hand motion scaling, which can improve the safety and precision of surgery, require a teleoperation control framework. This work presents a bimanual adaptive teleoperation (BMAT) control framework using SHER 2.0 and SHER 2.1 robotic systems. We integrate them with an adaptive force control (AFC) algorithm to automatically minimize the tool-sclera interaction force. The scleral forces are measured using two fiber Bragg grating (FBG)-based force-sensing tools. We compare the performance of the BMAT mode with a bimanual adaptive cooperative (BMAC) mode in a vessel-following experiment under a surgical microscope. Experimental results demonstrate the effectiveness of the proposed BMAT control framework in performing a safe bimanual telemanipulation of the eye without over-stretching it, even in the absence of registration between the two robots.
Index Terms:
Adaptive force control, bimanual adaptive teleoperation, steady-hand eye robot, fiber Bragg grating sensor, force-sensing surgical instrument.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I INTRODUCTION
Retinal vein occlusion (RVO) is the second most prevalent retinal vascular disease, impacting an estimated 28 million individuals globally, according to a meta-analysis study [1]. RVO occurs due to the occlusion of a retinal vein, resulting in eye problems and vision loss. There is no established standard for directly treating RVO with a surgical approach. Retinal vein cannulation (RVC) is a proposed solution to treat RVC by delivering a therapeutic agent directly into the affected vein [2]. Freehand RVC is highly challenging due to the micron scale of retinal veins and human hand tremor [3]. Of note, the diameter of the largest retinal veins is on the order of $150\mu$m [4], and the root mean square (RMS) amplitude of the hand tremor for an ophthalmic surgeon is reported as $182\mu$m [5].
@@
Several surgical robotic systems have been developed to filter out physiological hand tremors and improve the safety of retinal surgery. Some examples of these systems include the Steady Hand Eye Robot (SHER) [6, 7], and Preceyes [8], to name a few [9, 10, 11].
@@
Figure 1: Bimanual teleoperation architecture with adaptive force control algorithm. The system consists of two Phantom Omni robots, the SHER 2.0 and SHER 2.1, a surgical microscope, two FBG-based force-sensing surgical instruments attached to the SHERs’ end-effector to measure the tip force and sclera force, and two foot-pedals to activate and control the SHERs impedance. The figure shows the high-level adaptive force controller, a mid-level optimizer, and a low-level Galil joint velocity controller for SHER 2.0 and SHER 2.1.
@@
The SHER 2.0 and SHER 2.1 are two robot versions developed at Johns Hopkins University. The main control algorithm implemented on SHERs is an admittance-based control where the user and the robot cooperatively manipulate a surgical tool. Using a force/torque sensor attached to the robot handle, the exerted force by the user is measured and converted to a desired end-effector velocity to control the robot [7]. An important limitation of this control algorithm is that the force/torque sensor is not sensitive enough to measure the minimal interaction force existing on the surgical instrument at the sclerotomy point, which could result in applying excessive force to the sclera distorting and potentially causing damage. To overcome this issue, Ebrahimi et al. proposed an adaptive sclera force control algorithm [12] using an FBG-based force-sensing tool to automatically minimize the sclera force and keep it within a suggested safe range of 120 mN [13]. Notably, FBG sensors have applications in flexible needles or continuum robots, serving purposes such as shape sensing [14] or force sensing [15, 16]. In that study [12], a comparison of the adaptive force control algorithm and the admittance-based control algorithm performance was conducted using a vessel-following experiment in an eye phantom. A primary force-sensing surgical tool was attached to the SHER 2.1 and cooperatively manipulated by the user’s dominant hand. Simultaneously, a conventional surgical instrument was directly held by the non-dominant hand, serving as a secondary tool for re-orienting the eye under a surgical microscope [12]. Despite reducing the applied force to the sclera by the dominant hand compared to the admittance-based control mode, controlling the force magnitude on the secondary hand-held tool was impossible, and applying excessively unsafe force to the sclera could not be avoided. Also, this evaluation occurred within a unimanual cooperative control mode without teleoperation capability. Integrating this adaptive force control algorithm in a teleoperation modality has the potential to offer surgeons advanced capabilities, contributing to an improvement in patient safety. Esfandiari et al. implemented this adaptive force control algorithm in a teleoperation modality and compared its performance with a cooperative control mode [17]. However, these two studies, [12, 17], were conducted in a unimanual robotic framework, which is not a realistic strategy in real retinal surgeries, as they are typically performed bimanually. Doing a bimanual, as opposed to unimanual, robotic manipulation is a challenging problem and needs more safety measures such as task constraints [18, 19], robot registration [20], and camera calibration [19]. These techniques are typically used to dynamically track eye/head movement and satisfy geometrical constraints such as the remote center of motion (RCM) and the relative position of the two robots and the eye, to avoid stretching or compressing the eye. However, the accuracy of these registration-based methods may be adversely affected by eye or head movement, particularly in robotic systems utilizing external optical coherence tomography (OCT) [21].
To address these limitations, we locally minimize each robot’s tool-sclera interaction forces automatically and independently. At the same time, the surgeon bimanually manipulates the eye robots, either using two haptic interfaces in the teleoperation modality or through a hand-over-hand strategy in a cooperative modality. Of note, despite recent advancements in robot intelligence and autonomy, teleoperation is still a dominant method for controlling robots, particularly for performing complex tasks in highly unpredictable situations [22].
The contributions of this work are as follows:
  • We developed, for the first time in robot-assisted retinal microsurgery research applications, a bimanual adaptive teleoperation (BMAT) control framework integrated with an adaptive sclera force control algorithm on two retinal surgery robotic systems equipped with two FBG-based force-sensing surgical instruments. This framework enables surgeons to safely conduct bimanual teleoperation of the eye, preventing excessive stretch or compression of the sclera, even in the absence of registration of the two robots.
@@
  • We compared the performance of the proposed BMAT control mode with a bimanual adaptive cooperative control mode integrated with the same adaptive force control algorithm. This evaluation occurred within a vessel-following experiment on an eye phantom under a surgical microscope for both sitting and standing postures for each control mode. The objective was to assess and compare the performance of the BMAT and BMAC control modes and also the effect of the user’s posture on the performance of the proposed bimanual robotic framework.
Although the sitting posture is preferred in robot-assisted retinal microsurgery, we also study the performance of the proposed BMAT control mode in the standing posture to evaluate its safety for other potential applications in which surgeons may need to (partly) stand during surgery, for example, microvascular anastomosis [23].
@@
The remaining sections of this paper are organized as follows: Methods are represented in Section II. Section III describes the experimental setup and procedure. The experimental results are analyzed in Section IV and discussed in Section V. Section VI concludes the paper.
Figure 2: A diagram of the FBG-based force-sensing surgical instruments connected to the eye robots. These instruments measure the scleral force, tip forces, and insertion depth. The body coordinate, ${B}$, is attached to the robot end-effector, and the spatial coordinate system, ${S}$, is fixed to the robot base.
@@ II Methods
The objective of this study is to extend a previously developed adaptive teleoperation control mode [17] to two eye robots, the SHER 2.0 and SHER 2.1, and to assess the performance of the two control frameworks: bimanual adaptive cooperative mode and bimanual adaptive teleoperation mode (see Fig. 1), in both sitting and standing postures. In both control modes, the sclera force magnitude is automatically controlled (without the user’s commands) by an adaptive force control algorithm to avoid stretching and damaging the sclera. Following the suggestion of our clinical lead, we conducted a ”vessel-following” experiment, a prototypical task in retinal surgery [24]. The experimental setup prepared to perform the vessel-following experiment provides a measurement and recording of SHERs’ kinematic information, the user’s hand force/torque during interaction with the SHERs’ handle, and the force applied to the surgical tool at the sclera point and the tooltip.
II-A Robot Kinematics
The Steady-Hand Eye Robot is a 5-degree-of-freedom (5-DoF) robotic manipulator, comprising three translational motions along $XYZ$ cartesian coordinate followed by two rotational, yaw, and pitch degrees of freedom (see Fig. 1). It can carry out a surgical instrument, such as a cannulation tool or a microforceps, at its end-effector. There is no control over the roll motion around the instrument’s axis.
The kinematics of the SHER is described using two coordinate frames: the spatial coordinate $\{S\}$, which remains fixed in space and is situated at the robot’s base, and the body coordinate $\{B\}$, which is firmly attached to the robot’s handle and moves accordingly (see Fig. 2) [17].
Utilizing the homogeneous representation $g_{SB}:\Theta\rightarrow SE(3)$, the SHER end-effector coordinate can be described for any set of joint variables $\Theta\in\mathbb{R}^{5}$. Here, $g_{SB}\in\mathbb{R}^{4\times 4}$, which constitutes a matrix within the special Euclidean group $SE(3)$, represents the relative configuration of the body coordinate $\{B\}$ to the spatial coordinate $\{S\}$. Using the product of the exponential formula [25], the forward kinematics can be represented as follows
$g_{SB}(\Theta)=e^{\hat{\xi}_{1}\theta_{1}}...e^{\hat{\xi}_{5}\theta_{5}}g_{SB}(0),$ (1)
where $\theta_{i}$ represents the $i^{th}$ element of the joint variables $\Theta$, $\hat{\xi}i\in se(3)$ stands for the $i^{th}$ twist, and $g_{SB}(0)\in\mathbb{R}^{4\times 4}$ is the relative configuration of $\{B\}$ to $\{S\}$ when the robot is at home position ($\Theta=\textbf{0}$).
The twist $\hat{\xi}_{i}$ is calculated for the three translational joints ($i=1,2,3$) and the two rotational joints ($i=4,5$) using the following formula
$\hat{\xi}_{i}=\begin{cases}\quad\begin{bmatrix}0_{3\times 3}&v_{i}\\ 0_{1\times 3}&0\\ \end{bmatrix},\quad\quad\quad\quad\textrm{for}\quad i=1,2,3\vspace{0.2cm}\\ \quad\begin{bmatrix}\hat{\omega}_{i}&-\omega_{i}\times q_{i}\\ 0_{1\times 3}&0\\ \end{bmatrix},\quad\textrm{for}\quad i=4,5\end{cases}$ (2)
where $v_{i}$ is a unit vector representing the $i^{th}$ prismatic joint and is along with its positive direction, $\omega_{i}$ is a unit vector representing the $i^{th}$ revolute joint and is directed to the positive rotation axis (counterclockwise) of this joint, and $q_{i}$ denotes an arbitrary point on this rotation axis. Both $v_{i}$ and $\omega_{i}$ are represented in the $\{S\}$ coordinate when the robot is at home position.
Substituting the twist (2) into (1), it is possible to develop a mapping between the robot joint velocity vector $\dot{\Theta}\in\mathbb{R}^{5}$ and the robot end-effector velocity vector $V_{SB}^{b}\in\mathbb{R}^{6}$. Here, the superscript $b$ indicates that the velocity of the body coordinate $\{B\}$ relative to the spatial coordinate $\{S\}$ is expressed in $\{B\}$. This relationship is given by:
$V_{SB}^{b}=J_{SB}(\Theta)\dot{\Theta}$ (3)
where $J_{SB}\in\mathbb{R}^{6\times 5}$ is the robot Jacobian and is formulated as:
$J_{SB}(\Theta)$ $=\begin{bmatrix}\xi_{1}^{\dagger}&\dots&\xi_{5}^{\dagger}\end{bmatrix},$ (4)
$\xi_{i}^{\dagger}$ $=Ad^{-1}_{e^{\hat{\xi}_{i}\theta_{i}}\cdots e^{\hat{\xi}_{5}\theta_{5}}g_{SB}(0)}\xi_{i}$
where $Ad_{g}:\mathbb{R}^{6}\rightarrow\mathbb{R}^{6}$ represents the adjoint transformation of the rigid body transformation $g=(p,R)\in SE(3)$, and is defined as follows [25]:
$Ad_{g}=\begin{bmatrix}R&\hat{p}R\\ 0&R\end{bmatrix}_{6\times 6}.$ (5)
Figure 3: Control diagram of the bimanual teleoperation modality integrated with adaptive sclera force control algorithm.
II-B Adaptive Force Control of the SHER
This adaptive force control algorithm is automatically activated if each of the measured sclera force components, $F_{sx}$ and $F_{sy}$ (see Fig. 2), surpasses a specified safety threshold, $T_{s}$. Then, the robot is directed along a safe sclera force trajectory [12] minimizing the tool-sclera interaction force. The desired sclera forces, $f_{dx}$ and $f_{dy}$, are defined as exponentially decaying trajectories as follows
@@ II-B Adaptive Force Control of the SHER
[…]
@@
in which $\Delta{f}_{i}=F_{si}-f_{di}$ denotes the sclera force error, and $\Gamma_{i}$ and $K_{fi}$ represent constant gains for the adaptation law and the force tracking error, respectively. $\alpha_{i}$ stands for the estimated tissue compliance, and $\dot{f}_{di}$ is the derivative of $f_{di}$. This adaptive sclera force control algorithm applies to both cooperative and teleoperation control modalities. In the cooperative control mode, the remaining components of the SHER end-effector velocity are generated by the admittance-based controller. In contrast, in the teleoperation control mode, they are derived from the PHANTOM Omni’s end-effector velocity $\dot{X}_{o}$. The proposed BMAT control mode, integrated with an adaptive sclera force control algorithm, consists of three primary components: a high-level controller, a mid-level optimizer with joint constraints, and a low-level joint velocity controller (Galil), see Fig. 3.
The high-level controller consists of two primary components: Firstly, there is a kinematic controller that translates the end-effector velocity of the Omni, denoted as $\dot{X}_{o}$, into the SHER’s body coordinate as $\dot{X}_{o}^{b}$ (see Fig. 2). It is then pre-multiplied by $K=diag(\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4},\kappa_{5},\kappa_{6})$, a user-defined motion scaling matrix, to provide an intuitive motion on the SHER end-effector velocity, i.e., $\dot{X}_{des}^{b}=K\dot{X}_{o}^{b}$. Secondly, an adaptive sclera force controller that receives inputs from the sclera force components $F_{sx}$ and $F_{sy}$, which are measured in the body coordinate $\{B\}$ by the force-sensing instrument. This controller generates the desired adaptive end-effector velocity in the SHER body coordinate, denoted as $\dot{X}_{A}^{b}$ in (8).
This hybrid kinematic-force control method determines the desired SHER velocity by switching between $\dot{X}_{A}^{b}$ and $\dot{X}_{o}^{b}$ based on the magnitude of the sclera force. When the scleral force components fall below a safe threshold $T_{s}$, the SHER is controlled based on kinematic teleoperation using the Omni velocity $\dot{X}_{o}^{b}$; otherwise, the adaptive force controller takes charge, enforcing the SHER’s end-effector velocity through $\dot{X}_{A}^{b}$ to minimize scleral force components (see Fig. 3). As a result, the desired end-effector velocity of SHER, $\dot{X}_{des}^{b}$, is generated based on the following hybrid kinematic-force control algorithm:
@@
$\begin{bmatrix}(1-\Delta_{x})\kappa_{1}&0&\textbf{0}_{1\times 4}\\ 0&(1-\Delta_{y})\kappa_{2}&\textbf{0}_{1\times 4}\\ \textbf{0}_{4\times 2}&&diag(\kappa_{3},\kappa_{4},\kappa_{5},\kappa_{6})\end{bmatrix}\dot{X}_{o}^{b}$
@@
where $\Delta_{i}$ ($i=x$ or $y$) is a binary variable automatically set by the control switching policy (Algorithm 1) based on the magnitude of sclera force components $F_{si}$ ($i=x$ or $y$). $\Delta_{i}=1$ means the adaptive force control algorithm for the corresponding sclera force component ($x$ or $y$) is activated. Otherwise, when $\Delta_{i}=0$, the adaptive force control algorithm gets deactivated, and Equation (10) becomes equivalent to the kinematic control $\dot{X}_{des}^{b}=K\dot{X}_{o}^{b}$.
@@
       $t_{i}\leftarrow t$; $\triangleright$ Save current time $t$
       $\Delta_{i}\leftarrow 1$; $\triangleright$ Activate adaptive force control
       while  $|F_{si}|>0.75T_{s}$ do
             $\Delta_{i}\leftarrow 1$; $\triangleright$ Keep adaptive force control activated
             $f_{di}\leftarrow$ (6), (7); $\triangleright$ Calculate desired force
             $\dot{X}_{A_{i}}^{b}\leftarrow$ (8), (9); $\triangleright$ Calculate adaptive velocities
       end while
      $\Delta_{i}\leftarrow 0$; $\triangleright$ Switch to kinematic control for axis $i$
end if
Algorithm 1 Hybrid Kinematic-Force Control Switching Policy